a)     Suppose P (x, y) denotes the equation8, what will the truth values of the Propositions P (2,2), P (0,4).


1
Expert's answer
2022-01-10T15:24:52-0500

Let "P (x, y)" denotes the equation ""x+y=8"". Let us find the truth values of the Propositions "P (2,2), P (0,4)."

Since "2+2=4\ne 8," we conclude that the truth value of the Proposition "P (2,2)" is false.

Taking into account that "0+4=4\ne 8," we conclude that the truth value of the Propositions "P (0,4)" is false.